Rony has 2 more dimes in his pocket than Briana. The total value of their dimes is 2.40. How many dimes does each one have? Create an equation for this problem to find your answer
Answer by stanbon(75887)

Rony has 2 more dimes in his pocket than Briana. The total value of their dimes is 2.40. How many dimes does each one have? Create an equation for this problem to find your answer
---------------------
Equation:
R = B + 2
----
# of dimes = 240/10 = 24
------
Equation:
B+2 + B = 24
2B = 22
B = 11 (# of dimes Briana has)
B+2 = 13 (# of dimes Rony has)
-------------
Cheers,
Stan H.
-------------